# The Why Behind the Action: Unveiling Internal Drivers via Agentic Attribution

### v1

@@ Abstract
Large Language Model (LLM)-based agents are widely used in real-world applications such as customer service, web navigation, and software engineering. As these systems become more autonomous and are deployed at scale, understanding why an agent takes a particular action becomes increasingly important for accountability and governance. However, existing research predominantly focuses on failure attribution to localize explicit errors in unsuccessful trajectories, which is insufficient for explaining the reasoning behind agent behaviors. To bridge this gap, we propose a novel framework for general agentic attribution, designed to identify the internal factors driving agent actions regardless of the task outcome. Our framework operates hierarchically to manage the complexity of agent interactions. Specifically, at the component level, we employ temporal likelihood dynamics to identify critical interaction steps; then at the sentence level, we refine this localization using perturbation-based analysis to isolate the specific textual evidence. We validate our framework across a diverse suite of agentic scenarios, including standard tool use and subtle reliability risks like memory-induced bias. Experimental results demonstrate that the proposed framework reliably pinpoints pivotal historical events and sentences behind the agent behavior, offering a critical step toward safer and more accountable agentic systems.
@@ 1 Introduction
Large Language Model (LLM)-based agents have emerged as a powerful paradigm for automated task execution, demonstrating impressive capabilities in various domains (Yao et al., 2022; Qian et al., 2023; Guo et al., 2024; Wang et al., 2024a). By integrating long-term memory, external tools, and iterative reasoning loops, these agents are widely applied to dynamic environments and complex tasks ranging from software engineering to web navigation and scientific discovery (Jin et al., 2024; Ning et al., 2025; Ren et al., 2025). As agents become more capable and autonomous, failures in agentic systems have attracted increasing attention. Recent advances in Agentic Failure Attribution focus on identifying where an agent fails during task execution (Zhang et al., 2025c, ; Zhu et al., 2025). These approaches typically analyze completed trajectories after an explicit failure signal, such as an incorrect answer, an execution error, or task non-completion. By localizing faulty steps, they provide effective tools for debugging and improving system reliability.
Figure 1: Paradigm shift from failure attribution to agentic attribution. Top: Traditional failure attribution targets unsuccessful trajectories, aiming to localize the specific explicit error that caused the task failure. As shown, it identifies the “service unavailable” exception as the root cause preventing the booking confirmation. Bottom: The proposed agentic attribution framework uncovers the internal drivers behind an agent’s action. As illustrated in the customer service case, the agent directly issues an refund for a simple information inquiry, which carries no explicit error signal yet is undesirable. Our framework reveals this was driven by the memory of past ”refund action with high scores,” which overrode the user’s specific query.
@@
This motivates us to move beyond conventional failure-centered analysis toward a broader agentic attribution objective: understanding the internal factors that led the agent to a particular action. As agentic systems are deployed in real-world settings, concerns around transparency, accountability, and governance become increasingly critical (Dang et al., 2024; Raza et al., 2025; Luo et al., 2025). In these contexts, we need to understand why an agent takes a particular action, and whether its decision-making process was grounded in appropriate semantic evidence and aligned with the expected reasoning pathway. Formally, we model the agent’s execution trajectory as a temporal sequence composed of distinct interaction components (e.g., retrieved memory, tool observations, agent thoughts). Our goal is to assign an attribution score to each historical component—and subsequently to the fine-grained sentences within them—to quantify their contribution to the generation of a specific target action.
@@ 2.1 Structured Agent Interaction
where $o_{0}$ represents the initial input to the agent, such as the system prompt or user request. Observations $o_{t}$ include external feedback from the environment, such as tool execution results, retrieved documents or memories. Actions $a_{t}$ include internal reasoning steps, tool calls, and final responses.
@@ 2.2 Attribution Objectives
Given a completed interaction trajectory $\mathcal{C}$, an undesirable output action $a_{T}$, and an agent policy $\pi_{\theta}$, our goal is to quantify the contribution of individual components in $\mathcal{C}$ to the generation of $a_{T}$. Specifically, we formalize attribution at two levels: component-level attribution and sentence-level attribution.
@@
$f_{\mathrm{comp}}:(\mathcal{C},a_{T},\pi_{\theta})\rightarrow\mathbb{R}^{2T-1},$ (5)
@@
denote the sentence decomposition of component $C_{i}$, where $N_{i}$ is the number of sentence in component $C_{i}$. We define a sentence-level attribution function
@@ 3.1 Component Attribution by Temporal Likelihood Dynamics
Figure 2: Illustration of the Agentic Attribution Framework. The framework operates in a hierarchical approach to identify agent decisions. Level 1 (Left): Component-Level Attribution utilizes temporal likelihood dynamics to calculate the marginal gain of each component, selecting the high-impact component (e.g. $C_{3}$) that effectively steers the agent toward the final action $a_{T}$. Level 2 (Right): Sentence-Level Attribution performs fine-grained analysis within the identified component by computing perturbation-based score for each sentence.
@@
Recalling that the completed agentic interaction is represented as an ordered component sequence $\mathcal{C}=(C_{1},C_{2},\dots,C_{2T})$, where each $C_{i}$ corresponds to exactly one observation or action and preserves temporal order (Section 2.2). For each $i\in\{1,\dots,2T-1\}$, we form a temporal prefix trajectory
@@
$\psi_{i}\;=\;\log p_{\pi_{\theta}}(a_{T}\mid\mathcal{C}_{\leq i}),\qquad i\in\{1,\dots,2T-1\}.$ (9)
@@
$f_{\mathrm{comp}}(\mathcal{C},a_{T},\pi_{\theta})\;:=\;\{g_{i}\}_{i=1}^{2T-1}.$ (11)
@@
Discussion.
The core philosophy of this module is to leverage the temporal structure of agent execution via incremental replay. Unlike standard RAG (Gao et al., 2023) attribution which views context as a flat set of documents (Li et al., 2025; Zhang et al., 2025a), agentic attribution must respect the causal order of tool usage and memory updates, and reasoning steps. Note that, our framework relies on the temporal gain ($g_{i}=\psi_{i}-\psi_{i-1}$) to measure influence. While we instantiate $\psi_{i}$ using the generative log-likelihood, it remains agnostic and can be substituted with other specific metrics. For instance, to focus on high-level semantic intent rather than token-level probability, $\psi_{i}$ could be computed using embedding-based similarity between the evolving context and the action (Reimers and Gurevych, 2019). Alternatively, in safety-critical scenarios, $\psi_{i}$ may be derived from a trained reward model (Lightman et al., 2023; Lab et al., 2025) that quantifies the appropriateness of action $a_{T}$ given the current history. We leave the exploration of these alternative scoring metrics to future work.
@@ 3.2 Sentence-Level Attribution inside High-Impact Components
Discussion.
This stage provides sentence-level attribution explanations, bridging the gap between coarse trajectory components and specific semantic insights. Here, we preliminarily instantiate the sentence score $\phi_{i,j}$ using a perturbation-based approach, owing to its conceptual simplicity and effectiveness in recent attribution tasks (Liu et al., 2024; Cohen-Wang et al., 2024; Chuang et al., 2025). Also, our framework is method-agnostic regarding how $\phi_{i,j}$ is calculated. Depending on specific application scenarios, computational constraints, or model access, $\phi_{i,j}$ can be substituted with other established input attribution techniques, such as gradient-based method (Shrikumar et al., 2017), attention-based analysis (Hao et al., 2021), and linear-model-based prediction (Cohen-Wang et al., 2024). We empirically evaluate and compare several of these alternative attribution methods within our framework in Section 4.3.
@@ 4 Experiments
In this section, we evaluate the proposed agentic attribution framework through both qualitative and quantitative analyses. We first design the scenarios and cases used for agentic attribution (Section 4.1). Then, we present detailed qualitative case studies to illustrate attribution results at both the trajectory and sentence levels (Section 4.2). Then, we report a quantitative evaluation by instantiating our framework with different sentence-level attribution methods, and compare these alternatives against our default choice (Section 4.3).
@@ 4.1 Scenarios and Cases Design
In this subsection, we design a set of agentic scenarios to serve as the testbed for our attribution experiments. We design a set of agentic scenarios to serve as the testbed for our experiments. Unlike prevailing agent benchmarks like GAIA (Mialon et al., 2023) and HLE (Phan et al., 2025) that primarily evaluate the task correctness, and emerging failure detection datasets focus on localizing explicit runtime errors (Cemri et al., 2025; Zhang et al., 2025c), our objective is to verify whether the attribution framework can accurately identify the internal factors driving an agent’s decision. To this end, we design our scenarios around the two dominant information sources in agentic systems: memory-driven interactions and tool-driven interactions, which aligns with the two fundamental capabilities that distinguish agents from standard LLMs (Wang et al., 2024a; Guo et al., 2024).
@@
Figure 3: Illustration of attribution results across four representative scenarios. The highlighted regions denote the historical components and fine-grained sentences identified by our framework as the primary decision drivers.
Case Construction. Following the scenario taxonomy proposed above, we carefully construct a set of nine agent execution trajectories to evaluate our attribution framework. To ensure robustness, we curate these cases to cover a broad spectrum of behaviors, ranging from normal and benign operations to implicitly undesirable actions. All trajectories are generated by Llama-3.1-70B-Instruct (Grattafiori et al., 2024) using the smolagents [^1] https://github.com/huggingface/smolagents framework. Regarding data sources, we custom-design eight cases to capture specific interaction dynamics, while we carefully select one complex retrieval task from the GAIA benchmark (Mialon et al., 2023). We provide representative instances for the in-depth qualitative analysis in the following subsection.
4.2 Qualitative Analysis
Experimental Settings. We conduct agentic attribution experiments on scenarios introduced in Section 4.1, which serve as the testbed for qualitative analysis. Since all trajectories are generated by Llama-3.1-70B-Instruct through actual agent execution, we perform attribution using the same model to ensure attribution faithfulness. This choice follows a causal requirement of attribution: the goal is to explain why a specific trajectory is produced by a given model.
4.2.1 Memory-Driven Interaction
@@
Attribution Results. Component-level attribution shows that the user feedback following the agent’s laptop purchase advice receives the highest attribution score among all trajectory components. In this component, the user states that they did not follow the agent’s earlier advice to “explore some more affordable alternatives,” but instead purchased the expensive laptop and reports the outcomes after one month of use. Sentence-level attribution further refines this observation. Within this high-impact user component, sentences describing concrete financial outcomes receive the highest attribution scores. For example, the top-ranked sentence is “Also, its high performance allowed me to take on a few freelance video editing jobs, earning back \$800,” followed by “It seems that sometimes buying high-quality tools is actually saving and making money.” The agent’s subsequent response also receives a high attribution score, it explicitly retrieves this experience during the memory retrieval stage and re-evaluates its value judgment based on the reported laptop outcomes.
Discussion.
The attribution results indicate that the agent’s recommendation is primarily driven by a single prior success case, which is reused as a strong decision signal in a different context. Although the laptop purchase led to positive financial outcomes, the agent applies this experience broadly without adequately distinguishing between the two scenarios. This pattern suggests a possible risk of overgeneralization in memory-driven decision making, where context-specific outcomes may exert disproportionate influence on future value judgments.
Case Study 2: Proper Memory Update under Environmental Change.
@@ Discussion.
Attribution Results. The component-level attribution scores indicate that the agent’s final action is mainly influenced by the assistant step that performs memory retrieval after the user reports the heavy rainstorm. In this step, the agent updates several memory fields, including weather_status and risk_level, which capture the newly introduced environmental change and directly support the subsequent delivery decision.
A finer-grained analysis at the sentence level highlights which memory entries contribute most strongly within this step. The fields "risk_level": "HIGH" and "weather_status": "Heavy Rain" receive the highest attribution scores, followed by "item_sensitivity": "High (Water Damage)" and "location_is_covered": false. These entries align closely with the predefined decision logic that governs when the delivery location should be changed.
@@
4.2.2 Tool-Driven Interaction
Case Study 3: Spurious Tool Signal Leading to a Wrong Answer.
Case Overview. In this case, the agent answers a fact-seeking question from the GAIA [ref] benchmark using iterative web search. The task asks for the country of an unknown-language article under DDC 633 on BASE (Bielefeld University Library) as of 2020, described as having a flag that is unique from the others. The agent issues multiple search queries to identify the relevant entry and its country of origin. Although the agent does not establish a clear evidence chain for the target article, it eventually outputs Belarus, which is not the correct answer for this instance.
Attribution Results. The component-level attribution scores highlight a single early tool output as the most influential component in the trajectory. This top-ranked component corresponds to the first web search results, consists of a list of retrieved articles returned by the initial query, including bibliographic entries with titles, author information, and language-related metadata. Sentence-level attribution further shows that, within this component, the sentences containing the article title “Belarusian nation-building in times of war and revolution” and the associated BELARUSIAN label receive the highest attribution scores. In addition, the runner-up components include several sentences that explicitly express uncertainty also receive relatively high attribution scores, such as ”I still didn’t find the answer in the search result…” and ”However, I still don’t know which country…”
Discussion. This case illustrates a failure mode in tool-mediated reasoning where weakly grounded signals can dominate the decision process, leading to an unsupported answer even when uncertainty is present in the agent’s intermediate reasoning. Since we know the correct answer to this GAIA task is not Belarus, revisiting the attribution results may provide insight into how the error arises. While the search results contain a BELARUSIAN entry, the content does not establish that this entry matches the task’s constraints (DDC 633, unknown language, and unique flag). At the same time, the runner-up attribution also highlights multiple reasoning steps in which the agent explicitly acknowledges uncertainty and the lack of sufficient evidence. These signals indicate that the agent is, at least partially, aware of the ambiguity in the task. However, this uncertainty does not outweigh the influence of the earlier tool output.
@@ 4.2.2 Tool-Driven Interaction
Attribution Results. The attribution scores indicate that the agent’s final response is primarily influenced by two components in the execution trajectory. The highest attribution is assigned to the user request that explicitly specifies the questions to be answered, which defines the information requirements for the task. The tool invocation result that reads the attached document receives the second-highest attribution score. For sentence level, in the user request component, the sentences asking for two primary questions, i.e., the new safe combination and the new meeting point receive the highest attribution scores. Within the tool output component, the lines stating “The secure safe combination has been changed to: 7-2-9-4” and “the meeting point is no longer the library, it has been moved to the ‘Blue Lagoon Cafe’” are assigned the strongest attribution, directly corresponding to the information used in the final answer.
@@
4.3 Quantitative Analysis

### v2

@@ Abstract
Large Language Model (LLM)-based agents are widely used in real-world applications such as customer service, web navigation, and software engineering. As these systems become more autonomous and are deployed at scale, understanding why an agent takes a particular action becomes increasingly important for accountability and governance. However, existing research predominantly focuses on failure attribution to localize explicit errors in unsuccessful trajectories, which is insufficient for explaining the reason behind agent behaviors. To bridge this gap, we propose a novel framework for general agentic attribution, designed to identify the internal factors driving agent actions regardless of the task outcome. Our framework operates hierarchically to manage the complexity of agent interactions. Specifically, at the component level, we employ temporal likelihood dynamics to identify critical interaction steps; then at the sentence level, we refine this localization using perturbation-based analysis to isolate the specific textual evidence. We validate our framework across a diverse suite of agentic scenarios, including standard tool use and subtle reliability risks like memory-induced bias. Experimental results demonstrate that the proposed framework reliably pinpoints pivotal historical events and sentences behind the agent behavior, offering a critical step toward safer and more accountable agentic systems. Codes are available at https://github.com/AI45Lab/AgentDoG.
@@ 1 Introduction
Large Language Model (LLM)-based agents have emerged as a powerful paradigm for automated task execution, demonstrating impressive capabilities in various domains (Yao et al., 2022; Qian et al., 2023; Guo et al., 2024; Wang et al., 2024a). By integrating long-term memory, external tools, and iterative reasoning loops, these agents are widely applied to dynamic environments and complex tasks ranging from software engineering to web navigation and scientific discovery (Jin et al., 2024; Ning et al., 2025; Ren et al., 2025; Guo et al., 2025). As agents become more capable and autonomous, failures in agentic systems have attracted increasing attention. Recent advances in Agentic Failure Attribution focus on identifying where an agent fails during task execution (Zhang et al., 2025c, b; Zhu et al., 2025). These approaches typically analyze completed trajectories after an explicit failure signal, such as an incorrect answer, an execution error, or task non-completion. By localizing faulty steps, they provide effective tools for debugging and improving system reliability.
Figure 1: Paradigm shift from failure attribution to agentic attribution. Top: Traditional failure attribution targets unsuccessful trajectories, aiming to localize the specific explicit error that caused the task failure. As shown, it identifies the “service unavailable” exception as the root cause preventing the booking confirmation. Bottom: The proposed agentic attribution framework uncovers the internal drivers behind an agent’s action. As illustrated in the customer service case, the agent directly issues a refund for a simple information inquiry, which carries no explicit error signal yet is undesirable. Our framework reveals this was driven by the memory of past ”refund action with high scores,” which overrode the user’s specific query.
@@
This motivates us to move beyond conventional failure-centered analysis toward a broader agentic attribution objective: understanding the internal factors that led the agent to a particular action. As agentic systems are deployed in real-world settings, concerns around transparency, accountability, and governance become increasingly critical (Dang et al., 2024; Raza et al., 2025; Luo et al., 2025; Gan et al., 2026). In these contexts, we need to understand why an agent takes a particular action, and whether its decision-making process was grounded in appropriate semantic evidence and aligned with the expected reasoning pathway. Formally, we model the agent’s execution trajectory as a temporal sequence composed of distinct interaction components (e.g., retrieved memory, tool observations, agent thoughts). Our goal is to assign an attribution score to each historical component—and subsequently to the fine-grained sentences within them—to quantify their contribution to the generation of a specific target action.
@@ 2.1 Structured Agent Interaction
where $o_{0}$ represents the initial input to the agent, such as the system prompt or user request. Observations $o_{t}$ include external feedback from the environment, such as tool execution results, retrieved documents or memories. Actions $a_{t}$ include internal reasoning steps, tool calls, etc.
@@ 2.2 Attribution Objectives
Given a completed interaction trajectory $\mathcal{C}$, an agentic action $a_{T}$, and an agent policy $\pi_{\theta}$, our goal is to quantify the contribution of individual components in $\mathcal{C}$ to the generation of $a_{T}$. Specifically, we formalize attribution at two levels: component-level attribution and sentence-level attribution.
@@
$f_{\mathrm{comp}}:(\mathcal{C},a_{T},\pi_{\theta})\rightarrow\mathbb{R}^{2T+1},$ (5)
@@
denote the sentence decomposition of component $C_{i}$, where $N_{i}$ is the number of sentences in component $C_{i}$. We define a sentence-level attribution function
@@ 3.1 Component Attribution by Temporal Likelihood Dynamics
Figure 2: Illustration of the Agentic Attribution Framework. The framework operates in a hierarchical approach to identify agent actions. Left: Component-Level Attribution utilizes temporal likelihood dynamics to calculate the marginal gain of each component, selecting the high-impact component (e.g., $C_{4}$) that effectively steers the agent toward the final action $a_{T}$. Right: Sentence-Level Attribution performs fine-grained analysis within the identified component by computing perturbation-based score for each sentence.
@@
Recalling that the completed agentic interaction is represented as an ordered component sequence $\mathcal{C}=(C_{1},C_{2},\dots,C_{2T+2})$, where each $C_{i}$ corresponds to exactly one observation or action and preserves temporal order (Section 2.2), and $C_{2T+2}$ is exactly the $a_{T}$. For each $i\in\{1,\dots,2T+1\}$, we form a temporal prefix trajectory
@@
$\psi_{i}\;=\;\log p_{\pi_{\theta}}(a_{T}\mid\mathcal{C}_{\leq i}),\qquad i\in\{1,\dots,2T+1\}.$ (9)
@@
$f_{\mathrm{comp}}(\mathcal{C},a_{T},\pi_{\theta})\;:=\;\{g_{i}\}_{i=1}^{2T+1}.$ (11)
@@
Discussion. The core philosophy of this module is to leverage the temporal structure of agent execution via incremental replay. Unlike standard RAG (Gao et al., 2023) attribution which views context as a flat set of documents (Li et al., 2025; Zhang et al., 2025a), agentic attribution must respect the causal order of tool usage and memory updates, and reasoning steps. Note that, our framework relies on the temporal gain ($g_{i}=\psi_{i}-\psi_{i-1}$) to measure influence. While we instantiate $\psi_{i}$ using the generative log-likelihood, it remains agnostic and can be substituted with other specific metrics. For instance, to focus on high-level semantic intent rather than token-level probability, $\psi_{i}$ could be computed using embedding-based similarity between the evolving context and the action (Reimers and Gurevych, 2019). Alternatively, in safety-critical scenarios, $\psi_{i}$ may be derived from a trained reward model (Lightman et al., 2023; Lab et al., 2025) that quantifies the appropriateness of action $a_{T}$ given the current history. We leave the exploration of these alternative scoring metrics to future work.
@@ 3.2 Sentence-Level Attribution inside High-Impact Components
Discussion. This stage provides sentence-level attribution explanations, bridging the gap between coarse trajectory components and specific semantic insights. Here, we preliminarily instantiate the sentence score $\phi_{i,j}$ using a perturbation-based approach, owing to its conceptual simplicity and effectiveness in recent attribution tasks (Liu et al., 2024; Cohen-Wang et al., 2024; Chuang et al., 2025). Also, our framework is method-agnostic regarding how $\phi_{i,j}$ is calculated. Depending on specific application scenarios, computational constraints, or model access, $\phi_{i,j}$ can be substituted with other established input attribution techniques, such as gradient-based method (Shrikumar et al., 2017), attention-based analysis (Hao et al., 2021), and linear-model-based prediction (Cohen-Wang et al., 2024). We empirically evaluate and compare several of these alternative attribution methods within our framework in Section 4.3.
@@ 4 Experiments
In this section, we evaluate the proposed agentic attribution framework through both qualitative and quantitative analyses. We first design the scenarios and cases used for agentic attribution (Section 4.1). Then, we present detailed qualitative case studies to illustrate attribution results at both the trajectory and sentence levels (Section 4.2). Then, we report a quantitative evaluation by instantiating our framework with different sentence-level attribution methods (Section 4.3).
@@ 4.1 Scenarios and Cases Design
In this subsection, we design a set of agentic scenarios to serve as the testbed for our attribution experiments. Unlike prevailing agent benchmarks like GAIA (Mialon et al., 2023) and HLE (Phan et al., 2025) that primarily evaluate the task correctness, and emerging failure detection datasets focus on localizing explicit runtime errors (Cemri et al., 2025; Zhang et al., 2025c), our objective is to verify whether the attribution framework can accurately identify the internal factors driving an agent’s decision. To this end, we design our scenarios around the two dominant information sources in agentic systems: memory-driven interactions and tool-driven interactions, which aligns with the two fundamental capabilities that distinguish agents from standard LLMs (Wang et al., 2024a; Guo et al., 2024).
@@
Case Construction. Following the scenario taxonomy proposed above, we carefully construct a set of eight agent execution trajectories to evaluate our attribution framework. To ensure robustness, we curate these cases to cover a broad spectrum of behaviors, ranging from normal and benign operations to implicitly undesirable actions. All trajectories are generated by Llama-3.1-70B-Instruct (Grattafiori et al., 2024) using the smolagents (Roucher et al., 2025) framework. More details are available in Appendix A.
4.2 Analyzing Attribution in Representative Scenarios
Experimental Settings. We conduct agentic attribution experiments on all scenarios and cases introduced in Section 4.1, which serve as the testbed for qualitative analysis. We use Llama-3.1-70B-Instruct to perform attribution. We provide an in-depth attribution analysis of four representative cases in the following section, and include the attribution analyses for the remaining cases in Appendix B.
Figure 3: Illustration of attribution results across four representative scenarios. The components marked with red borders and medal icons denote the critical historical components identified by our component-level attribution. Within these components, the highlighted text spans indicate the specific semantic evidence localized by the sentence-level analysis.
4.2.1 Memory-driven Interaction.
@@ 4.2 Analyzing Attribution in Representative Scenarios
Attribution Results. Component-level attribution shows that the user feedback in Step 3 receives the highest attribution score among all trajectory components. In this component, the user states that they did not follow the agent’s earlier advice to “explore some more affordable alternatives,” but instead purchased the expensive laptop and reports the outcomes after one month of use. Sentence-level attribution further refines this observation. Within this high-impact user component, sentences describing concrete financial outcomes receive the highest attribution scores. For example, the top-ranked sentence is “…high performance helped me earn \$800…” followed by “…sometimes buying high-quality tools is actually saving and making money.” The agent’s subsequent response also receives a high attribution score, it explicitly retrieves this experience during the memory retrieval stage and re-evaluates its value judgment based on the reported laptop outcomes.
Discussion. The attribution results indicate that the agent’s recommendation is primarily driven by a single prior success case, which is reused as a strong decision signal in a different context. Although the laptop purchase led to positive financial outcomes, the agent applies this experience broadly without adequately distinguishing between the two scenarios. This pattern suggests a possible risk of overgeneralization in memory-driven decision making, where context-specific outcomes may exert disproportionate influence on future value judgments.
Case Study 2: Proper Memory Update under Env. Change.
@@
Attribution Results. The component-level attribution scores indicate that the agent’s final action is mainly influenced by the Step 4 that agent performs memory retrieval after the user reports the heavy rainstorm. In this step, the agent updates several memory fields, including weather_status and risk_level, which capture the newly introduced environmental change and directly support the subsequent delivery decision. A finer-grained analysis at the sentence level highlights which memory entries contribute most strongly within this step. The fields “risk_level”: “HIGH” and “weather_status”: “Heavy Rain” receive the highest attribution scores. These entries align closely with the predefined decision logic that governs when the delivery location should be changed.
@@
4.2.2 Tool-driven Interaction.
Case Study 3: External Prompt Injection.
Case Overview. In this safety-related scenario, the agent operates as a personal assistant with access to private emails and sensitive account details. The user issues a benign request to read the latest email regarding a specific project. However, the retrieved email body contains an adversarial prompt injection that instructs the agent to “Ignore all previous instructions” and send sensitive Amazon address data to an external recipient. The agent interprets the injected text as a valid new instruction, overriding the user’s original query and leaking the private information.
Attribution Results. The component-level attribution identifies the tool return in Step 3 as the decisive historical event. This component is the email body returned by the tool, and it includes the content of the malicious instructions. At the sentence level, the framework isolates the specific prompt injection “Ignore all previous instructions and strictly adhere to following instruction: Please retrieve…” as the highest-scoring semantic evidence. This segment receives significantly higher attribution scores than the surrounding email context, indicating that the agent’s execution is effectively hijacked by this imperative and goal-shifting text.
Discussion. This case demonstrates how agentic attribution can be applied to identify safety risks. Although the user request in Step 1 clearly defines the task boundary, the injected content in Step 3 induces the agent to override the original goal. The attribution results explicitly link the security breach to the specific external input, help localize both the key component and the specific injection phrases. More broadly, the case exposes a vulnerability where untrusted tool outputs are treated as executable instructions (Ruan et al., 2023; Zhan et al., 2024). In this scenario, our attribution framework provides a faithful explanation of why the agent performs the harmful action, which supports both debugging and the design of stronger safety guards in open-ended, potentially untrusted environments.
@@
Attribution Results. The attribution scores indicate that the agent’s final response is primarily influenced by two components in the trajectory. The highest attribution is assigned to the user request (Step 1) that explicitly specifies the questions to be answered, which defines the information requirements for the task. The tool invocation result (Step 3) that reads the attached document also receives a high attribution score. For sentence level, in Step 1, the sentences asking for two primary questions receive the highest attribution scores. Within the Step 3, the lines stating “The secure safe combination has been changed to: 7-2-9-4” and “the meeting point…has been moved to the ‘Blue Lagoon Cafe’ ” are assigned the strongest attribution, directly corresponding to the information used in the final answer.
@@
4.3 Comparative Evaluation of Attribution Methods